is x squared + y squared - 12x + 18y - 4 = 0 an ellipse, parabola, hyperbola or circle?
Answer by lwsshak3(11628)

is x squared + y squared - 12x + 18y - 4 = 0 an ellipse, parabola, hyperbola or circle?
**
x^2+y^2-12x+18y-4=0
Completing the squares
(x^2-12x+36)+(y^2+18y+81)=4+36+81=121
(x-6)^2+(y+9)^2=121
This is an equation of a circle with center at (6, -9) and radius=√121=11